A ball is kicked horizontally at 8 m/s from a cliff 80 meters high. Determine the velocity of the ball when it reaches the ground.



1
Expert's answer
2021-05-07T08:12:04-0400
"v^2=u^2+2gh\\v^2=8^2+2(9.8)(80)\\v=40.4\frac{m}{s}"